Illustration and solution required- please use the dorsal portion.



12-14. Two like charges of charge +10 nC are placed at the two corners of an equilateral triangle of sice 3cm. What is the magnitude of the total potential at the third corner?



15-17. A point charge q = 2.00μC is located at the origin. Find the electric potential due to this charge at point x = 4.0 m.

1
Expert's answer
2022-04-19T16:32:36-0400

1.

"\varphi_0=\frac{kq}r=3~kV,"

"\varphi=\sqrt{\varphi_0^2+\varphi_0^2+2\varphi_0^2\cos60°}=\varphi_0\sqrt 3=5.2~kV."

2.

"\varphi=\frac{kq}r=4.5~kV."